Gas fills a cylinder fitted with a frictionless piston. The initial pressure and volume are 40 MPa

and 0.05 dm3 respectively. The gas expands reversibly and polytropically to 0.5 MPa and 1 dm3

respectively. Calculate the index of expansion and the work done


1
Expert's answer
2021-09-07T14:07:35-0400

"p_1V_1^n=p_2V_2^n,"

"\frac{p_1}{p_2}=(\frac{V_2}{V_1})^n,"

"\frac{40}{0.5}=(\frac 1{0.05})^n,"

"80=20^n,"

"n\u22481.46,"


"A=\frac1{n-1}(p_1V_1-p_2V_2)=1025~J."